# ROADWork: A Dataset and Benchmark for Learning to Recognize, Observe, Analyze and Drive Through Work Zones

### v2

@@ 4 Observing Work Zones
As discussed in Section 3 we employ methods like Detic [115], OpenSeeD [105] and OMG-LLaVA [112]. We fine-tune instance segmentation models [33, 51] on ROADWork with fine-grained vocabulary for comparison. Our observations are in Table 5.All the foundation models [115, 111, 112] are dismal in segmenting these signs while Mask DINO [51] performs considerably better (+28.9 $AP$). Following Section 3, we adopt a similar strategy to exploit ROADWork videos and to propagate ground-truth labels using SAM2 [79] and obtain an additional improvement of +2.8 $AP$. However, overall performance of Mask R-CNN was still low (17.6 $AP$) compared to Mask DINO. We use simple copy-paste augmentation [28] which improves rare category segmentation. It further improved the accuracy of our Mask R-CNN [33] model by +7.1 $AP$.
$AP$ $AP50$ $AP75$ $AP_{s}$ $AP_{m}$ $AP_{l}$
@@
OpenSEED (COCO+O365) [111] 1.0 2.0 0.9 0.5 1.1 2.1
Detic (LVIS+I21K) [115] 1.1 1.5 1.3 0.3 1.8 2.2
OMG-LLaVA (Many Datasets) [112] 0 0.1 0.1 0 0 0.2
@@
Mask R-CNN [33] 17.6 26.2 19.5 11.3 21.2 36.0
Mask R-CNN w/ Copy Paste [28] 25.0 34.6 27.8 11.0 26.9 46.8
Mask DINO [51] 30.0 40.9 33.4 20.6 34.4 50.1
Mask DINO [51] w/ Video Prop. [79] 32.8 43.4 36.4 18.7 32.7 60.2
Table 5: Segmenting Fine-Grained Signs. We expand the “TTC sign” category to classify different kinds of TTC signs into separate categories. Open vocabulary methods [105, 115] fail to detect most of these objects with this expanded vocabulary. Supervised Mask DINO [51] improves considerably (+23.9 $AP$). Additionally, simple copy-paste [111] ensures simpler Mask R-CNN [33] baseline compares favorably with SOTA transformers.
@@
Gemini 2.5 Flash [16] 22.7 4.5 94.1 14.6 25.2
GPT-4o [39] 40.3 11.9 84.7 27.3 41.3
@@
Table 6: Reading Sign Text. The pretrained text spotting method Glass [81] performs well on nearby signs, but it fails to detect and read distant text (See Figure 7). Closed foundation models like GPT-4o [39] and Gemini 2.5 [16] are worse. A simple crop-rescale strategy composing a work zone detector with Glass [81] improves text spotting and recognition by +14.2% (1 - NED).
@@
Figure 7: Reading Sign Text. Glass [81] reads nearby signs but fails for distant signs (a-c). Dirt, occlusions and vandalism are issues for close-by signs too. In (d), prediction is “Keep Bight” .
@@
Figure 8: Generating Work Zone Descriptions. We present examples of work zone descriptions generated by a pretrained LLaVA [60] and a LoRA [36] fine-tuned LLaVA (with and without additional object context). The generated descriptions can be classified as factual, incorrect or uninformative. As we can observe, while pretrained LLaVA correctly identifies certain scene elements, it also generates many uninformative or inaccurate sentences. Fine-tuning LLaVA without object context reduces uninformative outputs but introduces hallucinated scene elements that do not exist. Incorporating LLaVA with our workzone detector to provide object context significantly improves faithfulness; however, it still misses some scene elements and spatial interactions (See San Antonio example’s ground truth).
@@
Reading Sign Text. ROADWork contains 360 different TTC sign texts. Sign texts are unconstrained and provide navigational instructions. Spotting and reading them accurately is important. We consider foundation models like GPT-4o [39] and Gemini 2.5 Flash [16] and a popular scene text spotting method, Glass [81]. We evaluate the spotter on TTC signs, and ignore other text in the scene. Foundation Models are unable to localize the text regions accurately, which is consistent with their underperformance on standard tasks [78]. We then try Glass [81], trained on the TextOCR [87] dataset. Glass accurately spots and reads most TTC signs, but fails on small signs far away (see Figure 7). Our annotations do not include character-level annotations required by Glass [81] so we did not fine-tune it on our data.
@@
We report sentence, word, and character level metrics for text spotting [81, 14]. From Table 6, we observe that while the precision of Glass [81] is high, the recall is low. Thus, we adopted a simple cropping strategy, composing Glass [81] with a coarse vocabulary work zone object detector [33] (Section 3). We infer the text spotter only on TTC sign crops. Our strategy significantly increased recall (+13.1%) at the expense of precision. Increasing the scale of the crops improved all metrics, such as word accuracy (+4.5%). There is a large performance gap for small signs – around 20% to 30% lower. The crop rescale strategy significantly improves word accuracy (+21.3%) for such signs.
@@
Summary. The ROADWork dataset enables studying foundation models to interpret sign graphics and spot sign text. Addressing failures, simple strategies such as copy-paste [28], video label propagation [79] and crop-rescaling image regions all improve downstream performance.
@@ 5 Analyzing Work Zones
Generating Work Zone Descriptions. Vision-Language Models (VLM) have shown strong performance in scene understanding in various scenarios [77, 52, 59, 60]. However, work zones are underrepresented in their training data. Our experiments show that while pre-trained BLIP [52] and LLaVA [60, 50] fail to generate informative work zone descriptions, fine-tuning on ROADWork enables them to produce accurate and contextually relevant descriptions.
@@
BLIP-361M [52] 0.2 4.3 8.6 1.6 3.9
LLaVA-1.5-7B [60] 0.4 11.0 9.4 0 9.9
GPT-4o [39] 0.9 16.4 14.4 0 -
Gemini 2.5 Flash [16] 0.6 13.2 11.7 0 -
@@
BLIP-361M [52] 20.7 21.5 43.7 83.5 41.3
LLaVA-1.5-7B [60] 27.0 24.7 48.0 112.1 42.7
@@
[60] + Pred. Objs 31.1 27.7 50.9 140.3 46.6
[60] + GT Objs 32.5 28.5 52.5 166.0 49.9
Table 7: Generating Work Zone Descriptions. Pretrained VLMs, BLIP [52] and LLaVA-1.5-7B [60], do not generate informative descriptions about work zones. Finetuning with LoRA [36] on ROADWork dramatically improves performance. LLaVA-1.5-7B [60] significantly outperforms BLIP [52], though larger LLaVA models perform similarly (See Appendix  C.2). Using object context enhances description faithfulness, but a performance gap remains between ground truth and uncertain object predictions.
Table 7 shows improvements with fine-tuning. LLaVA-7B [60] outperforms BLIP-361M [52] on all metrics by a large margin (e.g. +32.8 SPICE) – effectively utilizing the learned priors of a bigger model. Apart from fine-tuning, we added work zone object context by appending detection predictions from a trained work zone object detector, which further improves performance (+3.9 SPICE). Using ground truth objects as context establishes an upper bound, and a gap exists compared to providing uncertain object predictions (-3.3 SPICE). Figure 8 shows some generated descriptions from pre-trained LLaVA [60], with untrue and irrelevant sentences. In comparison, training on ROADWork dramatically improves description quality, and providing object context reduces hallucinations.
Summary. VLMs fail to describe work zones without the ROADWork dataset. Beyond fine-tuning, providing object context significantly improves performance. Appendix C.2 shows that larger foundation models, e.g. LLaVA-NEXT [50], also perform poorly, highlighting the need for the ROADWork dataset. Fine-tuning VLMs on our dataset preserves the overall performance and even improves COCO-Captions [10] (4.7 $\rightarrow$ 12.6 BLEU@4).
@@ 6 Driving Through Work Zones
Pathway Prediction In Work Zone Images. Work zones demand long-term planning beyond bird’s eye view (BEV) [15] representations, which are common for driving scenarios [90, 70, 37, 18] such as highways. Rare work zone objects and their spatial interactions are difficult to capture using BEV alone. Additionally, navigational goals in work zones are inherently multi-modal and cannot be effectively constrained within a fixed temporal window. Thus, we represent goals and pathways as spatial probability distributions in the image space, similar to long-horizon human path planning [63]. Perception cost maps for BEV planners [62, 24, 84] can be derived from these distributions.
@@ C R C R C R C R
$\theta={$$}$ 0.5  ° 43.7 53.6 31.2 38.4 67.2 75.3 67.1 75.4
$\theta={$$}$ 1  ° 63.5 73.8 53.8 60.5 82.8 87.2 82.7 87.0
$\theta={$$}$ 2  ° 79.0 85.9 73.8 83.0 90.5 93.0 90.4 93.0
$\theta={$$}$ 5  ° 89.6 93.2 91.4 96.8 96.0 97.3 96.0 97.3
$\theta={$$}$ 10  ° 94.5 97.0 95.7 98.9 98.5 99.3 98.5 99.3
Table 8: Pathway Prediction In Work Zones. YNet [63] models are trained on ROADWork dataset pathway data using pretrained segmentation models from Cityscapes (C) and ROADWork (R). Our metric measures the percentage of predictions where the angular error (AE) between the predicted and ground-truth pathway falls below a given threshold, considering the image’s field of view ($\sim$50°). The model trained with ROADWork segmentations consistently outperforms the Cityscapes-trained model at all thresholds. We also report results on a subset of high-curvature paths, hypothesizing that navigating irregular work zone paths is more challenging. As expected, $AE\%<\theta$ of predicted goal is lower (-15.4%) for this subset at tight thresholds (0.5°0.5°0.5  ° to 2°2  °).
We employ YNet [63]. We pre-train the segmentation model that YNet employs using the Cityscapes [17] dataset which contains no work zone objects and with the ROADWork dataset. Both models are trained on pathways from our dataset. Pixel-level metrics [63] do not account for the field of view of images (See Appendix C). Thus, we employ Percentage of Paths within Angular Error threshold $\theta$ ($AE\%<\theta$) as our metric. Table 8 presents our results. Our model, employing work zone semantic segmentations is better at all $\theta$ thresholds, than the model employing Cityscapes semantic segmentations. For example, at threshold 0.5°0.5°0.5  °, our model improves AEfor goals by +9.9% and AEfor pathways by +8.1% (see Figure 9 for examples). We hypothesize navigating irregular work zone paths that deviate from a straight line is more difficult. Thus we also report results on subsets of paths thresholded at different curvatures. AEat 0.5°0.5°0.5  ° for goals on paths with high curvature is lower by -15.4% than all paths. Our dataset demonstrates that curved paths are indeed more difficult.
@@
Figure 9: Pathway Prediction in Work Zones. We show goal and trajectory heatmaps predicted by YNet [63]. The model inputs are the image and the observed pathway. We display the future pathway (both computed from driving videos). The top row shows the predicted goal heatmaps, and the bottom row shows the predicted pathway heatmaps conditioned on a sampled goal. The predicted goal heatmap (highlighted with an orange box) closely aligns with the ground-truth, and the pathway appears plausible.
@@ Appendix A Related Works
Long-Tail Scenarios in Autonomous Driving. Driving datasets have evolved from KITTI [27] to more diverse collections like BDD100K [107], nuScenes [5], Mapillary [74] and Cityscapes [17], incorporating advanced sensor suites [7, 89, 23]. However, these datasets provide limited representation of long-tailed scenarios such as work zones - for instance, nuScenes contains only 19 driven sequences with work zones [86] out of 1000 scenes. Commercial self-driving vehicle deployments, while impressive in common situations, also find it difficult to navigate work zones, see Figure 10 for some failure examples collected from social media.
Prior research on long-tailed driving scenarios has largely focused on scene understanding. Datasets like CODA [53] (with 1500 scenes containing long-tailed objects), WildDash [108, 109] (with global weather and lighting variations), SegmentMeIfYouCan [6], and BDD-Anomaly [34] focus almost exclusively on recognition, rather than holistically addressing perception and navigation in scenarios like work zones. Another well-studied long-tailed scenario is driving in adverse weather. Despite data collection challenges, specialized datasets exist for fog [3, 83], night [107, 82], and snow [3, 4], although these also primarily target recognition. Figure 11 illustrates why recognition alone is insufficient for self-driving in work zones.
Work zones are complex, dynamic environments requiring multi-level understanding, yet they’ve received little attention due to the challenges in data mining [56] and task formulation [86]. To our best knowledge, no large-scale public dataset has specifically addressed work zones before our contribution. While the MMI Open Dataset [41] provides raw videos collected for road inspection, we develop scenario taxonomies and annotated work zones to create the ROADWork Dataset.
@@
Work Zones in Autonomous Driving. Prior research has addressed isolated work zone edge cases. For example, [30] recognize safety barriers using a laser scanner while [29] attempt to determine which lane lines define a valid lane in work zones. Later works [68, 86] attempted to classify and localize work zones, while others updated HD maps [67, 76] with additional work zone information. Concurrent work [45] has proposed segmenting construction areas in videos to detect continuous zones from a distance. However, no prior work systematically categorizes work zones, formulates tasks, or curates data for autonomous driving in these environments.
Figure 11: Recognition is Not Enough for Navigating Work Zones. Work zones are dynamic and rare occurrences, thus it is challenging to navigate through them. Depicted is a work zone navigation sequence with sign text detected by Glass [81], work zone descriptions generated by fine-tuned LLaVA-1.5A [60] (incorrect description indicated in red) and car trajectory estimated via COLMAP [85]. Observe that initially the worker is holding a “STOP” sign, but later switches to a “SLOW” sign as the truck passes, indicating that the road is open for traversal by the ego-vehicle. This example shows mere object recognition is not enough for navigation; continuous fine grained scene observation and global scene analysis are both necessary.
Language and Navigation in Work Zones. Unseen scenarios, such as newly appearing work zones along a route, pose a major challenge for autonomous driving. Work zones are a classic example of navigation in open-ended driving scenes, requiring a higher level of semantic generalization. Linguistic representations can help generalization, enabling introspective explanations [44] that improve action predictions [106].
Recently, Vision-Language Models (VLMs) [59] (and Large Language Models (LLMs)) have been increasingly applied to scene understanding, demonstrating state-of-the-art generalization and reasoning capabilities. Recent efforts [100, 64, 92, 49] have leveraged these VLMs and LLMs to redefine scene understanding and subsequently, motion planning. Navigating work zones require both visuospatial and linguistic abilities. To address this, we propose a work zone description benchmark to aid global scene understanding in workzones.
For navigation in work zones, we argue that long horizon trajectory forecasting is essential, as traditional structural cues like lanes may be unreliable. Prior works [63, 55, 31] explored a related setting: long horizon human trajectory forecasting. Inspired by this line of work, we propose a new pathway prediction problem and baselines to address tackle this challenge.

### v3

@@ 4 Observing Work Zones
As discussed in Section 3 we employ methods like Detic [zhou2022detecting], OpenSeeD [xu2023open] and OMG-LLaVA [zhang2024omgllava]. We fine-tune instance segmentation models [he2017mask, li2023maskdino] on ROADWork with fine-grained vocabulary for comparison. Our observations are in Table 5.All the foundation models [zhou2022detecting, zhang2023simple, zhang2024omgllava] are dismal in segmenting these signs while Mask DINO [li2023maskdino] performs considerably better (+28.9 $AP$). Following Section 3, we adopt a similar strategy to exploit ROADWork videos and to propagate ground-truth labels using SAM2 [ravi2024sam2] and obtain an additional improvement of +2.8 $AP$. However, overall performance of Mask R-CNN was still low (17.6 $AP$) compared to Mask DINO. We use simple copy-paste augmentation [ghiasi2021simple] which improves rare category segmentation. It further improved the accuracy of our Mask R-CNN [he2017mask] model by +7.1 $AP$.
APAP $AP50$ $AP75$ $AP_{s}$ $AP_{m}$ $AP_{l}$
@@
OpenSEED (COCO+O365) [zhang2023simple] 1.0 2.0 0.9 0.5 1.1 2.1
Detic (LVIS+I21K) [zhou2022detecting] 1.1 1.5 1.3 0.3 1.8 2.2
OMG-LLaVA (Many Datasets) [zhang2024omgllava] 0 0.1 0.1 0 0 0.2
@@
Mask R-CNN [he2017mask] 17.6 26.2 19.5 11.3 21.2 36.0
Mask R-CNN w/ Copy Paste [ghiasi2021simple] 25.0 34.6 27.8 11.0 26.9 46.8
Mask DINO [li2023maskdino] 30.0 40.9 33.4 20.6 34.4 50.1
Mask DINO [li2023maskdino] w/ Video Prop. [ravi2024sam2] 32.8 43.4 36.4 18.7 32.7 60.2
Table 5: Segmenting Fine-Grained Signs. We expand the “TTC sign” category to classify different kinds of TTC signs into separate categories. Open vocabulary methods [xu2023open, zhou2022detecting] fail to detect most of these objects with this expanded vocabulary. Supervised Mask DINO [li2023maskdino] improves considerably (+23.9 APAP). Additionally, simple copy-paste [zhang2023simple] ensures simpler Mask R-CNN [he2017mask] baseline compares favorably with SOTA transformers.
@@
Gemini 2.5 Flash [comanici2025gemini] 22.7 4.5 94.1 14.6 25.2
GPT-4o [hurst2024gpt4o] 40.3 11.9 84.7 27.3 41.3
@@
Table 6: Reading Sign Text. The pretrained text spotting method Glass [ronen2022glass] performs well on nearby signs, but it fails to detect and read distant text (See Figure 7). Closed foundation models like GPT-4o [hurst2024gpt4o] and Gemini 2.5 [comanici2025gemini] are worse. A simple crop-rescale strategy composing a work zone detector with Glass [ronen2022glass] improves text spotting and recognition by +14.2% (1 - NED).
@@
Figure 7: Reading Sign Text. Glass [ronen2022glass] reads nearby signs but fails for distant signs (a-c). Dirt, occlusions and vandalism are issues for close-by signs too. In (d), prediction is “Keep Bight” .
@@
Figure 8: Generating Work Zone Descriptions. We present examples of work zone descriptions generated by a pretrained LLaVA [liu2024improved] and a LoRA [hu2022lora] fine-tuned LLaVA (with and without additional object context). The generated descriptions can be classified as factual, incorrect or uninformative. As we can observe, while pretrained LLaVA correctly identifies certain scene elements, it also generates many uninformative or inaccurate sentences. Fine-tuning LLaVA without object context reduces uninformative outputs but introduces hallucinated scene elements that do not exist. Incorporating LLaVA with our workzone detector to provide object context significantly improves faithfulness; however, it still misses some scene elements and spatial interactions (See San Antonio example’s ground truth).
@@
Reading Sign Text. ROADWork contains 360 different TTC sign texts. Sign texts are unconstrained and provide navigational instructions. Spotting and reading them accurately is important. We consider foundation models like GPT-4o [hurst2024gpt4o] and Gemini 2.5 Flash [comanici2025gemini] and a popular scene text spotting method, Glass [ronen2022glass]. We evaluate the spotter on TTC signs, and ignore other text in the scene. Foundation Models are unable to localize the text regions accurately, which is consistent with their underperformance on standard tasks [ramachandran2025well]. We then try Glass [ronen2022glass], trained on the TextOCR [singh2021textocr] dataset. Glass accurately spots and reads most TTC signs, but fails on small signs far away (see Figure 7). Our annotations do not include character-level annotations required by Glass [ronen2022glass] so we did not fine-tune it on our data.
@@
We report sentence, word, and character level metrics for text spotting [ronen2022glass, chng2019icdar2019]. From Table 6, we observe that while the precision of Glass [ronen2022glass] is high, the recall is low. Thus, we adopted a simple cropping strategy, composing Glass [ronen2022glass] with a coarse vocabulary work zone object detector [he2017mask] (Section 3). We infer the text spotter only on TTC sign crops. Our strategy significantly increased recall (+13.1%) at the expense of precision. Increasing the scale of the crops improved all metrics, such as word accuracy (+4.5%). There is a large performance gap for small signs – around 20% to 30% lower. The crop rescale strategy significantly improves word accuracy (+21.3%) for such signs.
@@
Summary. The ROADWork dataset enables studying foundation models to interpret sign graphics and spot sign text. Addressing failures, simple strategies such as copy-paste [ghiasi2021simple], video label propagation [ravi2024sam2] and crop-rescaling image regions all improve downstream performance.
@@ 5 Analyzing Work Zones
Generating Work Zone Descriptions. Vision-Language Models (VLM) have shown strong performance in scene understanding in various scenarios [radford2021learning, li2022blip, liu2023llava, liu2024improved]. However, work zones are underrepresented in their training data. Our experiments show that while pre-trained BLIP [li2022blip] and LLaVA [liu2024improved, li2024llavanext-strong] fail to generate informative work zone descriptions, fine-tuning on ROADWork enables them to produce accurate and contextually relevant descriptions.
@@
BLIP-361M [li2022blip] 0.2 4.3 8.6 1.6 3.9
LLaVA-1.5-7B [liu2024improved] 0.4 11.0 9.4 0 9.9
GPT-4o [hurst2024gpt4o] 0.9 16.4 14.4 0 -
Gemini 2.5 Flash [comanici2025gemini] 0.6 13.2 11.7 0 -
@@
BLIP-361M [li2022blip] 20.7 21.5 43.7 83.5 41.3
LLaVA-1.5-7B [liu2024improved] 27.0 24.7 48.0 112.1 42.7
@@
[liu2024improved] + Pred. Objs 31.1 27.7 50.9 140.3 46.6
[liu2024improved] + GT Objs 32.5 28.5 52.5 166.0 49.9
Table 7: Generating Work Zone Descriptions. Pretrained VLMs, BLIP [li2022blip] and LLaVA-1.5-7B [liu2024improved], do not generate informative descriptions about work zones. Finetuning with LoRA [hu2022lora] on ROADWork dramatically improves performance. LLaVA-1.5-7B [liu2024improved] significantly outperforms BLIP [li2022blip], though larger LLaVA models perform similarly (See Appendix  C.2). Using object context enhances description faithfulness, but a performance gap remains between ground truth and uncertain object predictions.
Table 7 shows improvements with fine-tuning. LLaVA-7B [liu2024improved] outperforms BLIP-361M [li2022blip] on all metrics by a large margin (e.g. +32.8 SPICE) – effectively utilizing the learned priors of a bigger model. Apart from fine-tuning, we added work zone object context by appending detection predictions from a trained work zone object detector, which further improves performance (+3.9 SPICE). Using ground truth objects as context establishes an upper bound, and a gap exists compared to providing uncertain object predictions (-3.3 SPICE). Figure 8 shows some generated descriptions from pre-trained LLaVA [liu2024improved], with untrue and irrelevant sentences. In comparison, training on ROADWork dramatically improves description quality, and providing object context reduces hallucinations.
Summary. VLMs fail to describe work zones without the ROADWork dataset. Beyond fine-tuning, providing object context significantly improves performance. Appendix C.2 shows that larger foundation models, e.g. LLaVA-NEXT [li2024llavanext-strong], also perform poorly, highlighting the need for the ROADWork dataset. Fine-tuning VLMs on our dataset preserves the overall performance and even improves COCO-Captions [chen2015microsoft] (4.7 $\rightarrow$ 12.6 BLEU@4).
@@ 6 Driving Through Work Zones
Pathway Prediction In Work Zone Images. Work zones demand long-term planning beyond bird’s eye view (BEV) [claussmann2019review] representations, which are common for driving scenarios [teng2023motion, mcnaughton2011motion, hu2023planning, dauner2023parting] such as highways. Rare work zone objects and their spatial interactions are difficult to capture using BEV alone. Additionally, navigational goals in work zones are inherently multi-modal and cannot be effectively constrained within a fixed temporal window. Thus, we represent goals and pathways as spatial probability distributions in the image space, similar to long-horizon human path planning [mangalam2021goals]. Perception cost maps for BEV planners [lu2014layered, ferguson2008motion, salzmann2020trajectron] can be derived from these distributions.
@@ C R C R C R C R
[…]
Table 8: Pathway Prediction In Work Zones. YNet [mangalam2021goals] models are trained on ROADWork dataset pathway data using pretrained segmentation models from Cityscapes (C) and ROADWork (R). Our metric measures the percentage of predictions where the angular error (AE) between the predicted and ground-truth pathway falls below a given threshold, considering the image’s field of view ($\sim$50°). The model trained with ROADWork segmentations consistently outperforms the Cityscapes-trained model at all thresholds. We also report results on a subset of high-curvature paths, hypothesizing that navigating irregular work zone paths is more challenging. As expected, $AE\%<\theta$ of predicted goal is lower (-15.4%) for this subset at tight thresholds (0.5° to 2°).
We employ YNet [mangalam2021goals]. We pre-train the segmentation model that YNet employs using the Cityscapes [cordts2016cityscapes] dataset which contains no work zone objects and with the ROADWork dataset. Both models are trained on pathways from our dataset. Pixel-level metrics [mangalam2021goals] do not account for the field of view of images (See Appendix C). Thus, we employ Percentage of Paths within Angular Error threshold $\theta$ ($AE\%<\theta$) as our metric. Table 8 presents our results. Our model, employing work zone semantic segmentations is better at all $\theta$ thresholds, than the model employing Cityscapes semantic segmentations. For example, at threshold 0.5°, our model improves AE% for goals by +9.9% and AE% for pathways by +8.1% (see Figure 9 for examples). We hypothesize navigating irregular work zone paths that deviate from a straight line is more difficult. Thus we also report results on subsets of paths thresholded at different curvatures. AE% at 0.5° for goals on paths with high curvature is lower by -15.4% than all paths. Our dataset demonstrates that curved paths are indeed more difficult.
@@
Figure 9: Pathway Prediction in Work Zones. We show goal and trajectory heatmaps predicted by YNet [mangalam2021goals]. The model inputs are the image and the observed pathway. We display the future pathway (both computed from driving videos). The top row shows the predicted goal heatmaps, and the bottom row shows the predicted pathway heatmaps conditioned on a sampled goal. The predicted goal heatmap (highlighted with an orange box) closely aligns with the ground-truth, and the pathway appears plausible.
@@ Appendix A Related Works
Long-Tail Scenarios in Autonomous Driving. Driving datasets have evolved from KITTI [geiger2012we] to more diverse collections like BDD100K [yu2020bdd100k], nuScenes [nuscenes], Mapillary [neuhold2017mapillary] and Cityscapes [cordts2016cityscapes], incorporating advanced sensor suites [chang2019argoverse, Waymodataset, WaymoMotion]. However, these datasets provide limited representation of long-tailed scenarios such as work zones - for instance, nuScenes contains only 19 driven sequences with work zones [shi2021work] out of 1000 scenes. Commercial self-driving vehicle deployments, while impressive in common situations, also find it difficult to navigate work zones, see Figure 10 for some failure examples collected from social media.
Prior research on long-tailed driving scenarios has largely focused on scene understanding. Datasets like CODA [li2022coda] (with 1500 scenes containing long-tailed objects), WildDash [zendel2018wilddash, zendel2022wilddash2] (with global weather and lighting variations), SegmentMeIfYouCan [chan2021segmentmeifyoucan], and BDD-Anomaly [hendrycks2019bddanomaly] focus almost exclusively on recognition, rather than holistically addressing perception and navigation in scenarios like work zones. Another well-studied long-tailed scenario is driving in adverse weather. Despite data collection challenges, specialized datasets exist for fog [bijelic2020seeing, sakaridis2021acdc], night [yu2020bdd100k, sakaridis2019guided], and snow [bijelic2020seeing, burnett2023boreas], although these also primarily target recognition. Figure 11 illustrates why recognition alone is insufficient for self-driving in work zones.
Work zones are complex, dynamic environments requiring multi-level understanding, yet they’ve received little attention due to the challenges in data mining [liang2024aide] and task formulation [shi2021work]. To our best knowledge, no large-scale public dataset has specifically addressed work zones before our contribution. While the MMI Open Dataset [roadbotics-open] provides raw videos collected for road inspection, we develop scenario taxonomies and annotated work zones to create the ROADWork Dataset.
@@
Work Zones in Autonomous Driving. Prior research has addressed isolated work zone edge cases. For example, [gumpp2009recognition] recognize safety barriers using a laser scanner while [graf2012probabilistic] attempt to determine which lane lines define a valid lane in work zones. Later works [mathibela2013roadwork, shi2021work] attempted to classify and localize work zones, while others updated HD maps [mathibela2012can, pannen2020keep] with additional work zone information. Concurrent work [kim2024rosa] has proposed segmenting construction areas in videos to detect continuous zones from a distance. However, no prior work systematically categorizes work zones, formulates tasks, or curates data for autonomous driving in these environments.
Figure 11: Recognition is Not Enough for Navigating Work Zones. Work zones are dynamic and rare occurrences, thus it is challenging to navigate through them. Depicted is a work zone navigation sequence with sign text detected by Glass [ronen2022glass], work zone descriptions generated by fine-tuned LLaVA-1.5A [liu2024improved] (incorrect description indicated in red) and car trajectory estimated via COLMAP [schoenberger2016sfm]. Observe that initially the worker is holding a “STOP” sign, but later switches to a “SLOW” sign as the truck passes, indicating that the road is open for traversal by the ego-vehicle. This example shows mere object recognition is not enough for navigation; continuous fine grained scene observation and global scene analysis are both necessary.
Language and Navigation in Work Zones. Unseen scenarios, such as newly appearing work zones along a route, pose a major challenge for autonomous driving. Work zones are a classic example of navigation in open-ended driving scenes, requiring a higher level of semantic generalization. Linguistic representations can help generalization, enabling introspective explanations [kim2018textual] that improve action predictions [xu2020explainable].
Recently, Vision-Language Models (VLMs) [liu2023llava] (and Large Language Models (LLMs)) have been increasingly applied to scene understanding, demonstrating state-of-the-art generalization and reasoning capabilities. Recent efforts [wang2023drive, mao2023gpt, tian2024drivevlm, li2024llada] have leveraged these VLMs and LLMs to redefine scene understanding and subsequently, motion planning. Navigating work zones require both visuospatial and linguistic abilities. To address this, we propose a work zone description benchmark to aid global scene understanding in workzones.
For navigation in work zones, we argue that long horizon trajectory forecasting is essential, as traditional structural cues like lanes may be unreliable. Prior works [mangalam2021goals, liang2020garden, gupta2018social] explored a related setting: long horizon human trajectory forecasting. Inspired by this line of work, we propose a new pathway prediction problem and baselines to address tackle this challenge.